# Adversarially Perturbed Precision Matrix Estimation

### v1

@@ 6.2 Experiments with Synthetic Data
For each model, we run simulations by the QUIC algorithm (Hsieh et al., 2014) with dimension $d=20$ and sample sizes $n=\{30,50,70\}$. We consider an equally spaced grid for the perturbation magnitude $\delta$ on $[0.01,0.3]$ in problem (14) and select the regularization parameter $\lambda$ in problem (16) over an equally spaced grid on $[0.01,0.8]$. For each setting, we run two methods 50 times and record the average and standard error.
6.2.1 Sparsity Recovery and Accuracy
We first compare the performance, i.e., sparsity recovery ability and accuracy, based on Kullback–Leibler (KL) loss, TNR, and MCC. The KL loss is defined as follows:
$\mathrm{KL}=-\log|\bar{C}|+\mathrm{tr}(\bar{C}\Sigma)-\left(-\log|\Sigma^{-1}|+d\right),$
where $\Sigma$ is the ground-truth covariance matrix, and $\bar{C}$ is the estimation of the precision matrix.
We visualize the experimental results for the 6 models in Figure 1, Figure 2, and Figure 3. In the figures, “Perturbed” refers to the performance of our proposed method, whereas “Standard” indicates the performance of the standard $\ell_{1}$-regularized method. The x-axis shows the values of the $\delta$ and, $\lambda$ while the y-axis shows the different metrics.
According to the figures, we could conclude that the proposed method (14) has the ability of sparsity recovery since the TNR of the perturbed method can reach or be close to $1$. In addition, the performances of two methods are comparable in terms of accuracy and sparsity recovery ability.
Figure 1: Accuracy and Sparsity Recovery on $AR(1)$ and $AR(2)$
Figure 2: Accuracy and Sparsity Recovery on $AR(3)$ and $AR(4)$
Figure 3: Accuracy and Sparsity Recovery on the Circle model and the Star model
6.2.2 Adversarial Robustness
In addition to the sparsity recovery ability and the accuracy discussed previously, we also evaluate the adversarial robustness of the methods.
In each simulation trial, three independent datasets are generated from the true distribution: a training set $X_{\text{train}}$ of size $n$, a validation set of size $n$, and a test set $X_{\text{test}}$ of size $5n$. For each sample size $n$, we first identify the optimal choice of $\delta$ for the proposed method and $\lambda$ for the standard $\ell_{1}$-regularized method by minimizing the value of $\mathrm{KL}$ on the validation set. We then derive the estimation $\bar{C}_{1}$ from the training data. Next, we generate adversarial examples from the test data based on $\bar{C}_{1}$. The adversarial perturbation is constructed using a gradient-sign method defined as:
$X_{\text{adv}}=X_{\text{test}}+\varepsilon\text{sign}(X_{\text{test}}\bar{C}_{1}),$
where $\varepsilon$ controls the magnitude of the adversarial perturbation. We vary $\varepsilon$ across a grid from 0 to $1$.
Then, we run the two estimation methods based on the perturbed data $X_{\text{adv}}$ and obtain the estimation $\bar{C}_{2}$. The robustness of each method is evaluated by the $\mathrm{KL}$ loss of $\bar{C}_{2}$.
$\text{adversarial KL}=-\log|\bar{C}_{2}|+\mathrm{tr}(\bar{C}_{2}\Sigma)-\left(-\log|\Sigma^{-1}|+d\right)$
A smaller adversarial KL loss indicates that the estimator is more robust to adversarial data perturbation.
The results are visualized in Figure 4, Figure 5, and Figure 6. The x-axis shows the values of $\varepsilon$ while the value of the adversarial KL. The figures demonstrate that the proposed perturbed method is more adversarially robust than the standard $\ell_{1}$-regularized model for most of the scenarios.
Figure 4: Adversarial Robustness on $AR(1)$ and $AR(2)$
Figure 5: Adversarial Robustness on $AR(3)$ and $AR(4)$
Figure 6: Adversarial Robustness on the Circle model and the Star model
@@ 6.3 Experiments with Real Data
The dataset comprises 7129 gene expression profiles collected from 72 patients, of whom 47 were diagnosed with Acute Lymphoblastic Leukemia (ALL) and 25 with Acute Myeloid Leukemia (AML). We first select the top $d$ genes with the highest marginal variance. The dimension $d$ is varied across the set $\{20,30,40,50,60\}$. The data is centered and scaled to unit variance within the training set and test set, respectively. Following the Linear Discriminant Analysis (LDA) framework, the classification rule relies on the LDA score. For the observation $\bm{x}$, the LDA score is defined by
@@
The experiment employs a nested cross-validation scheme. In the outer loop, the data is split into 10 stratified folds. For each training run, we additionally carry out an inner stratified 5‑fold cross‑validation to tune the parameters ($\delta$ in the perturbed method (14) and $\lambda$ in the standard $\ell_{1}$-regularized method (16)) through a grid search within $[0.01,1]$, choosing the value that has the best classification accuracy. With the optimal parameter, we train the model on the full training set and then evaluate the model performance on the held‑out test fold. We replicate this whole scheme $50$ times.
To evaluate the performance, we use ACC, MCC, TNR, and TPR. The average value and the standard error of the metrics for 50 replications are recorded in Table 2. Notably, our perturbed method yields consistently higher ACC and MCC, at the cost of a slight decrease in TPR. In conclusion, the proposed method demonstrates superior efficacy in the precision matrix estimation for the leukemia dataset, outperforming the standard $\ell_{1}$-regularized method.
Method ACC MCC TNR TPR
$d=20$ Perturbed 0.868(0.00360) 0.691 (0.01009) 0.627 (0.01014) 0.995 (0.00137)
Standard 0.864 (0.00313) 0.679 (0.00882) 0.611 (0.00951) 0.997 (0.00101)
$d=30$ Perturbed 0.910 (0.00290) 0.792 (0.00806) 0.741(0.00886) 1.000 (0.000500)
Standard 0.902 (0.00301) 0.775 (0.00760) 0.716 (0.00871) 1.000 (0.000000)
$d=40$ Perturbed 0.933 (0.00285) 0.850 (0.00732) 0.809 (0.00840) 1.000 (0.00040)
Standard 0.921 (0.00305) 0.819 (0.00807) 0.771 (0.00928) 1.000 (0.00000)
$d=50$ Perturbed 0.927 (0.00312) 0.833 (0.00841) 0.798 (0.00926) 0.994 (0.00153)
Standard 0.916 (0.00268) 0.808 (0.00722) 0.760 (0.00778) 0.998 (0.000928)
$d=60$ Perturbed 0.929(0.00247) 0.839 (0.00712) 0.806 (0.00727) 0.993 (0.00143)
Standard 0.912 (0.00305) 0.802 (0.00779) 0.757 (0.00800) 0.994 (0.00148)
Table 2: Average Value and Standard Error of the Metrics
@@ 7 Discussion
This paper introduces a versatile perturbed estimation framework for the precision matrix and derives the associated regularization effects and asymptotic distributions. Among different perturbation geometries, the $\ell_{2}$- and $\ell_{\infty}$-perturbations are specifically analyzed. Notably, we demonstrate that the proposed framework recovers Wasserstein shrinkage estimation, achieves sparsity recovery, and improves adversarial robustness. In addition to these properties, future directions may include extending our framework to handle complex data structures, including the unknown group structures studied in (Cheng et al., 2025) and the compositional data investigated in (Zhang et al., 2025).
Appendix A Proofs
A.1 Proof of Theorem 2.1
The proof for Theorem 2.1 is based on the following lemma.
@@ Appendix A Proofs
$\mathbb{E}_{\mathbb{P}_{n}}\left[\|C\bm{x}\|_{1}\right]=\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{d}|[C]_{j}\bm{x}_{i}|$ $\leq\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{d}\sum_{k=1}^{d}|[C]_{kj}||[\bm{x}_{i}]_{k}|$
$=\sum_{j=1}^{d}\sum_{k=1}^{d}|[C]_{kj}|\left(\frac{1}{n}\sum_{i=1}^{n}|[\bm{x}_{i}]_{k}|\right).$

### v2

@@ 6.1 Experiments with Synthetic Data
For each model, we run simulations by the QUIC algorithm (Hsieh et al., 2014) with dimension $d=30$ and sample sizes $n=\{20,30,40\}$. We consider equally spaced grids for the perturbation magnitude $\delta$ in problem (9) and the regularization parameter $\lambda$ in problem (16) over an equally spaced grid on $[0.01,1]$. The value of $\delta$ and $\lambda$ is chosen based on the Bayesian Information Criterion. For each setting, we run two methods 100 times and record the average and standard error.
We report the results of the perturbed method and the $\ell_{1}$-regularized method in Tables 26. Since our primary interest lies in heteroscedastic settings, we additionally report the performance of the $\ell_{1}$-regularized method applied to standardized data. Across the majority of scenarios, the perturbed method attains higher ACC and MCC. This observation is consistent with our theoretical results on improved model selection consistency under heteroscedastic models.
$n$ Method ACC MCC TNR TPR
2020 Perturbed 0.876 (0.000) 0.232 (0.004) 0.994 (0.000) 0.095 (0.002)
$\ell_{1}$ 0.233 (0.001) 0.036 (0.004) 0.132 (0.001) 0.903 (0.004)
$\ell_{1}$ std 0.869 (0.000) 0.016 (0.005) 0.999 (0.000) 0.006 (0.002)
$30$ Perturbed 0.877 (0.000) 0.243 (0.003) 0.994 (0.000) 0.102 (0.002)
$\ell_{1}$ 0.698 (0.002) -0.002 (0.003) 0.769 (0.002) 0.229 (0.005)
$\ell_{1}$ std 0.871 (0.001) 0.051 (0.011) 0.998 (0.001) 0.026 (0.007)
$40$ Perturbed 0.879 (0.000) 0.263 (0.004) 0.994 (0.001) 0.115 (0.003)
$\ell_{1}$ 0.709 (0.001) 0.010 (0.004) 0.781 (0.002) 0.232 (0.005)
$\ell_{1}$ std 0.875 (0.001) 0.148 (0.018) 0.993 (0.002) 0.096 (0.014)
Table 2: Average Value and Standard Error of the Metrics for $AR(2)$
$n$ Method ACC MCC TNR TPR
2020 Perturbed 0.818 (0.000) 0.218 (0.003) 0.998 (0.000) 0.069 (0.002)
$\ell_{1}$ 0.318 (0.012) 0.000 (0.005) 0.203 (0.020) 0.796 (0.022)
$\ell_{1}$ std 0.807 (0.000) 0.005 (0.002) 1.000 (0.000) 0.001 (0.001)
$30$ Perturbed 0.820 (0.000) 0.228 (0.003) 0.998 (0.000) 0.074 (0.002)
$\ell_{1}$ 0.662 (0.001) -0.047 (0.002) 0.780 (0.002) 0.171 (0.002)
$\ell_{1}$ std 0.808 (0.000) 0.027 (0.007) 0.999 (0.000) 0.009 (0.003)
$40$ Perturbed 0.819 (0.000) 0.223 (0.003) 0.998 (0.000) 0.072 (0.002)
$\ell_{1}$ 0.672 (0.001) -0.035 (0.003) 0.791 (0.002) 0.173 (0.003)
$\ell_{1}$ std 0.813 (0.001) 0.085 (0.014) 0.996 (0.001) 0.045 (0.009)
Table 3: Average Value and Standard Error of the Metrics for $AR(3)$
$n$ Method ACC MCC TNR TPR
2020 Perturbed 0.761 (0.000) 0.203 (0.003) 0.999 (0.000) 0.058 (0.001)
$\ell_{1}$ 0.407 (0.014) -0.029 (0.005) 0.324 (0.029) 0.650 (0.032)
$\ell_{1}$ std 0.747 (0.000) 0.005 (0.002) 1.000 (0.000) 0.001 (0.000)
$30$ Perturbed 0.761 (0.000) 0.199 (0.002) 1.000 (0.000) 0.054 (0.001)
$\ell_{1}$ 0.618 (0.001) -0.078 (0.002) 0.776 (0.002) 0.151 (0.001)
$\ell_{1}$ std 0.748 (0.000) 0.012 (0.004) 1.000 (0.000) 0.002 (0.001)
$40$ Perturbed 0.760 (0.000) 0.195 (0.002) 1.000 (0.000) 0.053 (0.001)
$\ell_{1}$ 0.631 (0.001) -0.059 (0.002) 0.793 (0.002) 0.154 (0.001)
$\ell_{1}$ std 0.749 (0.001) 0.031 (0.007) 0.999 (0.001) 0.011 (0.004)
Table 4: Average Value and Standard Error of the Metrics for $AR(4)$
$n$ Method ACC MCC TNR TPR
2020 Perturbed 0.866 (0.008) 0.534 (0.016) 0.861 (0.008) 0.929 (0.009)
$\ell_{1}$ 0.837 (0.009) 0.496 (0.006) 0.826 (0.010) 0.989 (0.002)
$\ell_{1}$ std 0.641 (0.009) 0.157 (0.006) 0.639 (0.010) 0.666 (0.012)
$30$ Perturbed 0.886 (0.006) 0.581 (0.011) 0.879 (0.006) 0.981 (0.005)
$\ell_{1}$ 0.861 (0.001) 0.524 (0.002) 0.851 (0.001) 0.996 (0.001)
$\ell_{1}$ std 0.644 (0.003) 0.208 (0.004) 0.635 (0.003) 0.772 (0.006)
$40$ Perturbed 0.902 (0.004) 0.611 (0.007) 0.896 (0.004) 0.997 (0.001)
$\ell_{1}$ 0.873 (0.001) 0.546 (0.002) 0.864 (0.001) 0.999 (0.000)
$\ell_{1}$ std 0.628 (0.002) 0.236 (0.003) 0.612 (0.002) 0.854 (0.006)
Table 5: Average Value and Standard Error of the Metrics for the Star model
$n$ Method ACC MCC TNR TPR
2020 Perturbed 0.622 (0.014) 0.310 (0.008) 0.594 (0.015) 0.993 (0.002)
$\ell_{1}$ 0.398 (0.015) 0.181 (0.007) 0.355 (0.016) 0.980 (0.002)
$\ell_{1}$ std 0.633 (0.012) 0.320 (0.009) 0.606 (0.013) 0.995 (0.001)
$30$ Perturbed 0.706 (0.004) 0.363 (0.004) 0.684 (0.005) 1.000 (0.000)
$\ell_{1}$ 0.587 (0.005) 0.283 (0.003) 0.557 (0.005) 0.998 (0.001)
$\ell_{1}$ std 0.713 (0.011) 0.381 (0.008) 0.692 (0.012) 0.999 (0.000)
$40$ Perturbed 0.707 (0.003) 0.363 (0.003) 0.685 (0.004) 1.000 (0.000)
$\ell_{1}$ 0.606 (0.004) 0.294 (0.002) 0.577 (0.004) 1.000 (0.000)
$\ell_{1}$ std 0.666 (0.011) 0.346 (0.009) 0.642 (0.012) 1.000 (0.000)
Table 6: Average Value and Standard Error of the Metrics for the Circle model
@@ 6.2 Experiments with Real Data
The dataset comprises 7129 gene expression profiles collected from 72 patients, of whom 47 were diagnosed with Acute Lymphoblastic Leukemia (ALL) and 25 with Acute Myeloid Leukemia (AML). We first select the top $d$ genes with the highest marginal variance. The dimension $d$ is varied across the set $\{20,30,40,50,60\}$. The data is then centered and standardized. Following the Linear Discriminant Analysis (LDA) framework, the classification rule relies on the LDA score. For the observation $\bm{x}$, the LDA score is defined by
@@
The experiment employs a nested cross-validation scheme. In the outer loop, the data is split into 10 stratified folds. For each training run, we additionally carry out an inner stratified 5‑fold cross‑validation to tune the parameters, $\delta$ in the perturbed method (9) and $\lambda$ in the standard $\ell_{1}$-regularized method (16), through a grid search within $[0.01,1]$, choosing the value that has the best MCC. With the optimal parameter, we train the model on the full training set and then evaluate the model performance on the held‑out test fold. We replicate this whole scheme $100$ times.
To evaluate performance, we report ACC, MCC, TNR, and TPR. We treat AML as the positive class and ALL as the negative class, thus TPR corresponds to sensitivity for AML and TNR corresponds to specificity for ALL. The average metrics and their standard errors are computed across the 100 replications, where the standard error is calculated from the variability of the replicate-level means.
Notably, across the considered values of $d$, the perturbed method generally achieves higher ACC, MCC, and TPR. Overall, these results suggest that the proposed perturbed estimator can provide improved precision matrix estimates for this dataset, translating into better LDA classification performance compared with the $\ell_{1}$-regularized approach.
$d$ Method ACC MCC TNR TPR
2020 Perturbed 0.866 (0.002) 0.682 (0.005) 0.997 (0.001) 0.618 (0.005)
$\ell_{1}$ 0.863 (0.002) 0.670 (0.006) 0.997 (0.001) 0.607 (0.006)
$30$ Perturbed 0.921 (0.002) 0.822 (0.005) 1.000 (0.000) 0.771 (0.006)
$\ell_{1}$ 0.918 (0.002) 0.817 (0.004) 1.000 (0.000) 0.764 (0.005)
$40$ Perturbed 0.938 (0.002) 0.863 (0.004) 0.998 (0.001) 0.823 (0.004)
$\ell_{1}$ 0.937 (0.002) 0.860 (0.004) 0.998 (0.001) 0.820 (0.005)
$50$ Perturbed 0.944 (0.001) 0.878 (0.003) 0.997 (0.001) 0.844 (0.004)
$\ell_{1}$ 0.939 (0.001) 0.866 (0.004) 0.998 (0.001) 0.826 (0.004)
$60$ Perturbed 0.938 (0.001) 0.864 (0.004) 0.990 (0.001) 0.838 (0.004)
$\ell_{1}$ 0.935 (0.001) 0.859 (0.003) 0.993 (0.001) 0.825 (0.004)
Table 7: Average Value and Standard Error of the Metrics for leukemia dataset
@@ 7 Discussion
In addition to robustness and model selection consistency, future directions may include extending our framework to handle more complex data structures, including the unknown group structures studied in (Cheng et al., 2025) and the compositional data investigated in (Zhang et al., 2025).
While the adversarial perturbations investigated in our framework are sample-wise, as defined in (2), an intriguing direction for future research would be to extend this to feature-wise perturbations. In the linear regression, xu2008robust has shown that robust linear regression framework under the feature-wise perturbation is equivalent to LASSO. Inspired by this fact, the computational and statistical explorations of feature-wise perturbed precision matrix estimation is of much interest.
Appendix A Proof of Theorem 2.1
The proof for Theorem 2.1 is based on the following lemmas.
@@
$\mathbb{E}_{\mathbb{P}_{n}}\left[\|C\bm{x}\|_{1}\right]=\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{d}|[C]_{j}\bm{x}_{i}|$ $\leq\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{d}\sum_{k=1}^{d}|[C]_{kj}||\bm{x}^{i}_{k}|$
$=\sum_{j=1}^{d}\sum_{k=1}^{d}|[C]_{kj}|\left(\frac{1}{n}\sum_{i=1}^{n}|\bm{x}^{i}_{k}|\right).$